OpenStudy (anonymous):

5by^3/8a^3d-w^2/2ab^2

OpenStudy (anonymous):

no more trig?

OpenStudy (anonymous):

my assignments go back and forth because im on bed rest and I've missed so much

OpenStudy (anonymous):

common denominator is \[8a^3b^2d\] if i am reading the question correctly

OpenStudy (anonymous):

\[\frac{5by^3}{8a^3d}-\frac{w^2}{2ab^2}\]right?

OpenStudy (anonymous):

exactly.

OpenStudy (anonymous):

so you have to multiply the first one top and bottom by \[4b^2\] and the second one top and bottom by \[4a^2d\]

OpenStudy (anonymous):

this is a pain

OpenStudy (anonymous):

\[\frac{5by^3}{8a^3d}-\frac{w^2}{2ab^2}\] i am sorry i messed up

OpenStudy (anonymous):

yeah your telling me.. ive been doing this one assignment for 5 hours

OpenStudy (anonymous):

let me just write what you need to do

OpenStudy (anonymous):

okay:)

OpenStudy (anonymous):

\[\frac{5by^3b^2}{8a^3b^2d}-\frac{8a^2dw^2}{8a^3b^2d}\] \[

OpenStudy (anonymous):

\[\frac{5b^3y^3-8a^2d^2w^2}{8a^3b^2d}\]

OpenStudy (anonymous):

how the flipper does that even simplify with all the variables not being the same for the numerator?